OpenStudy (anonymous):

Consider the functions f(x) = 3x − 2 and g(x) = x2 − 1. What is the value of (f − g)(−2)?

OpenStudy (jdoe0001):

(f − g)(−2) => f( -2 ) - g( -2 )

OpenStudy (anonymous):

whats the answer though

OpenStudy (jdoe0001):

well, plug in the values and substract them :)

OpenStudy (anonymous):

can you explain it to me though???

OpenStudy (jdoe0001):

\(\bf f(x) = 3x − 2\\ g(x) = x^2 − 1\\ f(-2)= 3(-2) − 2\\ g(-2)= (-2)^2 − 1\)

OpenStudy (anonymous):

so next would be 4-1 and that's 3 so the answer is 3?

OpenStudy (jdoe0001):

$$ \bf f(-2)= \color{blue}{3(-2) − 2}\\ g(-2)= \color{green}{(-2)^2 − 1}\\ f(-2)-g(-2) = [\color{blue}{3(-2) − 2}] - [\color{green}{(-2)^2 − 1}] $$

OpenStudy (anonymous):

I am so lost....

OpenStudy (anonymous):

@Zale101

OpenStudy (anonymous):

I'm help less UGH

OpenStudy (jdoe0001):

$$\bf f(-2)-g(-2) = [\color{blue}{3(-2) − 2}] - [\color{green}{(-2)^2 − 1}]\\ f(-2)-g(-2) = [-6-2]-[4-1]$$

OpenStudy (jdoe0001):

so -6-2 = -8 -[4-1] => -[-3] => 3 -8+3 :)

OpenStudy (anonymous):

so -5?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

im sorry I suck at math

OpenStudy (jdoe0001):

keep in mind that when grouping - ( 2 + 3) => -1*2 + -1*3 => -2-3